OpenStudy (anonymous):

what is the solution to -19*^7sqrt15x+8=-38

OpenStudy (campbell_st):

so is the question 1. \[-19^7\times \sqrt{15x + 8} = -39\] or 2. \[-19 \times \sqrt[7]{15x + 8} = -38\]

OpenStudy (anonymous):

number "2"

OpenStudy (campbell_st):

ok... divide both sides by -19 so \[\sqrt[7]{15x + 8} = 2\] now raise both sides to the 7th power \[(\sqrt[7]{15x + 8})^7 = 2^7\] so \[15x + 8 = 128\] now solve for x

OpenStudy (anonymous):

8?

OpenStudy (campbell_st):

makes sense to me. well done